Mir Status Report 19 November 1997

Date: Wednesday, November 19, 1997// MD 106/55

Mir Station System/Subsystem Activities

Health of the major systems continue to remain unchanged (Kvant Electron working (periodically turned off/on for power), 11 gyrodynes holding attitude, Condensate System working, Urine Recycling System working).

After continuing discussions with different Russians on last Friday's solar array test, we have learned the experiment was performed to test the Cooperative Solar Array and NOT the newly installed Kvant solar array. Additionally, we learned that the disconnection of the Base Block arrays and the newly installed solar array were not part of this test.

This morning, Victor Blagov told us he attributes the experiment failure to an errant disconnection of one of the Cooperative Solar array connectors. Victor stated this experiment was performed to support the Contract (per Jim Nise, the Contract item is 3FA14). Victor also told us the experiment is scheduled to be repeated Thursday afternoon (after 16:00 DMT) to accomplish the contract requirement.

Today, we also received clarification on the sequence used to bring the gyrodynes back on-line. The Kvant-2 gyrodynes were spun down nominally since power to the module was not interrupted. The spinup of these gyrodynes occurred nominally. However, after the gyrodynes were integrated into the control loop, an out of limit signal was sensed by the precession unit which resulted in two gyrodynes (#2 and #6) spinning down. The out of limit signal was a result of excessive torques when controlling the Mir stack with 6 gyrodynes in one module. These two gyrodynes were spun up after one of the Kvant gyrodynes were back on line and continue to operate nominally.

The Kvant gyrodynes were not spun down nominally since their power is supplied from the Base Block. When testing the gyrodynes for start-up, the ground was unsuccessful in gaining an indication of health status. After two failed attempts to test the health of the gyrodynes, the ground team gave the crew a GO to spin up one gyrodyne (Kvant #4). The gyrodyne spun up nominally and the crew was directed to bring the remaining gyrodynes on-line. After discussing the situation with the motion control experts, the start-up test requires an inhibit command to be lifted prior to the testing. It is believed this inhibit was not released prior to the test.

Prior to the motion control failure, the crew had been working on the Electron unit in Kvant-2 which failed to run due to the O2 sensor in the H2 vent line sensing elevated pressures. The ground controllers believe the sensor may be clogged and have instructed the crew to remove and clean the sensor. These activities will be postponed until late this week or early next week.

Consumables used during the motion control recovery consist of 5 Oxygen canisters, 1 American LiOH canister, approximately 25 kg of Mir propellant, and approximately 10 kg of Soyuz prop.

In addition to the motion control recovery procedures, this week's systems work has focused on preparation and installation of the Base Block Vozdukh. The majority of the crew's time has been to clear out the space where the Vozdukh will be installed. The crew has had multiple questions concerning which hardware and cables were acceptable to move and are concerned that the allotted space will not be sufficient. Vozdukh activities are scheduled for 4-5 hrs/day for the rest of the week.

The ground is watching a decreasing efficiency on the air conditioner compressor. Since a spare is available on-board, the crew will be instructed to use the system as normal until the compressor fails. At that time, the compressor will be replaced with the on-board spare.

Tues evening tm data indicated the ShSO pressure at 278 mmHg.

At the last communication session of the day the crew reported problems with installing the Vozdukh unit in the Base Block. The crew said that there are some compatibility problems with existing pipes on board and that the specialist will have to be consulted tomorrow.

Mir Payload Activities

MG:BT:I:39 BEM/BSTC-M & BTR Facility Maintenance
FB:HSR:I:23 CHAPAT Experiment Computer Monitoring
HLS:BP:I:10 MIR Crew: Completion of Computerized Questionnaire Packet
HLS:BP:I:13 NASA Crew: Completion of Computerized Questionnaire Packet
OPS:FAC:I:11 Data Files Downlink
SMP:ENV:I:67 Micro Air and Surface Colony Count T+2
SMP:ENV:I:37 MCD Colony Count T+2
HLS:MET:I:101 Night Headband-Presleep *
HLS:MET:I:92 Night Headband-Postsleep *
*Cosmonaut Activities

Experiment Hardware -

IPCG - Currently the IPCG computer is processing Fluid System 2 Experiment 1. This was initiated Tuesday evening and is scheduled to run for 66 hours. At the completion of this run, Dave will be instructed to terminate System 2 due to the leakage of the cell. Fluid System 3 will not be initiated until after the next EVA. Dave will be sent instructions via radiogram on the proper procedure to terminate the automatic computer processing of the second system and reset the computer counter to prepare for System 3. Additionally, Dave will be instructed to allow pressure to decrease in the growth cell if, while injecting solution, he should encounter resistance and increased back pressure while turning the hand crank. Although the procedures call for the injection to occur over a sixty second period, there are indications that the injection process has taken up to five minutes on the ground.

BIO-3D - There are currently no hardware anomalies with either the BSTC or the BTR. After restoration of power to the experiment facilities following the power outage in the core module, all chamber temperatures of the BSTC and the temperature of the BTR are reading within nominal limits. SMP/Medical--A no medical impact PMC was completed today with Dave. Dave reported that he was not sure where the CCK was located on MIR. He was notified that it might be located under the refrigerator in the base block and we asked if he would confirm this for us. He confirmed that "there has been absolutely no medical problems as a result of the spill from IPCG."

Environmental Parameters (previous 24 hour period):
Parameter Min. Max. Nominal Range
P (mmHg) BB 688 670 660 - 850
ppO2 (mmHg) 149 153 140 - 200
ppCO2 (mmHg) 3.9 4.4 up to 8.0
ppH20 (mmHg) 13.4 14.7 8 - 14
Humidity (%) 49 51 30- 75
Temp (xC) BB 27.3 28.5 18 - 28

Priroda Temperature (deg C) at BV1 Fan: 27.3

Communications Summary / A/G with Consultant Group
09:54 - 09:58 IPCG, CHAPAT
16:00 - 16:06 CREAM, IPCG
20:34 - 20:36 Water sampling, IPCG

Packet -One English Form 24 file was sent today.

  • submit to reddit


Support SpaceRef, NASA Watch and the Astrobiology Web on Patreon.

Join our mailing list

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on November 19, 1997 9:01 PM.

Mir Status 20 October 1997 was the previous entry in this blog.

Mir Status Report 20 November 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.